OpenStudy (anonymous):

Pam’s monthly food budget is equal to 40% of her monthly rent payment. If her food budget is $200 a month, how much is her rent payment each month?

OpenStudy (anonymous):

set up a proportion: 40% = 40/100 and the other fraction would be 200/x. Your proportion is 40/100 = 200/x. Cross multiply and solve for x to get the answer.